OpenStudy (anonymous):

(3/x-3) + (x/x+3) = ? Show the work? I dont know how to make the denominators the same.

OpenStudy (anonymous):

You need to find the LCM (aka LCD) of each of the denominators

OpenStudy (anonymous):

(x-3)(x+3)?

OpenStudy (anonymous):

Yep. So now you want to multiply each fraction by the one that it's missing. The first fraction is missing the x+3. So multiply that fraction by (x+3)/(x+3)

OpenStudy (anonymous):

That will make the denominator match your LCD, and the numerator will be 3(x+3)

OpenStudy (anonymous):

Then for the second fraction, it's missing x-3, so multiply it by (x-3)/(x-3)

OpenStudy (anonymous):

That will again make the denominator match your LCD, and the numerator will be x(x-3)

OpenStudy (anonymous):

Now you can add the two fractions and simplify.

OpenStudy (anonymous):

What do you have after you did all that, just to make sure you didn't make a mistake

OpenStudy (anonymous):

Well, I got 3x.... & i dont think that sounds right...

OpenStudy (anonymous):

What did you have for the fractions when you changed denominators?

OpenStudy (anonymous):

\[\frac{3(x+3)}{(x+3)(x-3)} + \frac{x(x-3)}{(x+3)(x-3)}\] Right?

OpenStudy (anonymous):

You did that alot faster than I could... but yeah, thats what I got!

OpenStudy (anonymous):

Ok so now when you add them: \[=\frac{3(x+3) + x(x-3)}{(x+3)(x-3)}\]

OpenStudy (anonymous):

Then dont the (x+3)'s on the left side of the addition sign cancel out?

OpenStudy (anonymous):

They do, but you don't want to do that cause then your denominators will be different again

OpenStudy (anonymous):

And you won't be able to add them

OpenStudy (anonymous):

Shoot, okay... & where do you go from there.. just distribute on the top?

OpenStudy (anonymous):

Yep, distribute then combine like terms

OpenStudy (anonymous):

Let me know what you get in the numerator and I'll check it

OpenStudy (anonymous):

Isn't it 5x+6?

OpenStudy (anonymous):

wait.

OpenStudy (anonymous):

2x+9

OpenStudy (anonymous):

Sike; x^2 + 9.. took me long enough to make a thousand mistakes.

OpenStudy (anonymous):

\[3(x+3) + x(x-3)\]\[=3x + 9 + x^2 - 3x\]\[= x^2 + 9\]

OpenStudy (anonymous):

Right, haha, sorry.(:

OpenStudy (anonymous):

Right. So your answer is: \[\frac{x^2+9}{(x+3)(x-3)}\]

OpenStudy (anonymous):

What do you do with the bottom? Foil it?

OpenStudy (anonymous):

Sure if you want to

OpenStudy (anonymous):

Probably that's how they want it, though I prefer factored personally

OpenStudy (anonymous):

Right? Me too. Well thank you very much! You were lots of help!

OpenStudy (anonymous):

Certainly. Just remember when you need to add/subtract fractions with dissimilar denominators you must use this process: 1) factor the denominators 2) find the LCD for all of the fractions 3) multiply each fraction (top & bottom) by the factors it's missing 4) add/subtract the fractions together over the common denominator 5) distribute and simplify the numerator 6) cancel any common factors in the top/bottom to reduce

OpenStudy (anonymous):

Alright; will do! Thanks(: